OpenStudy (anonymous):

If i have a pizza that has a diameter of 16 inches and is cut into 12 slices what is the interior angle and the arc length of the slices, and the area of each slice?

OpenStudy (whpalmer4):

If the pizza is circular with a diameter of 16 inches, what is the circumference of the pizza? \[C = 2\pi r = \pi d\]where \(r\) is the radius and \(d\) is the diameter. If you cut the pizza into 12 identical slices, the round edge of each slice should occupy 1/12 of the circumference, no? That gives you the arc length. The interior angle will be 1/12 of 360 degrees. Finally, the area will be 1/12 of the area of the entire pizza. For a circle, \[A = \pi r^2 = \frac{1}{4}\pi d^2\]